The body temperatures in degrees Fahrenheit of a sample of adults in one small town are: 96.8 96.7 98.6 97.4 99.9 97.1 98.5 97.9 97 Assume body temperatures of adults are normally distributed. Based on this data, find the 95% confidence interval of the mean body temperature of adults in the town. Enter your answer as an open-interval (i.e., parentheses) accurate to 3 decimal places. Assume the data is from a normally distributed population.
Answer #1

From the given data,

Mean = 97.77

Standard deviation = 1.063

Degrees of freedom = n - 1 = 9 - 1 = 8

For 8 degrees of freedom and 95% confidence,

Critical t value = 2.306

Hence,

95% confidence interval will be:

(96.950, 98.584)